(30 points) Acme Engineering needs to buy a punch press. Two companies have submitted bids to supply Acme with a punch press. Johnsons Punches will charge $150,000 to deliver and install the press. They estimate that maintenance for the press will be $4,000 per year. They state that their press will save Acme Engineering $89,000 per year. Riser Machines will charge $205,000 to deliver and install their press. Riser estimates that that maintenance for their press will be $4,300 per year. They estimate that their press will save Acme $86,000 per year. The Machine from Acme should last for 5 years; the machine from Riser will last for 10 years. Both machines should have a salvage value of $15,000 at the end of their useful lives, whenever that occurs. Assuming an interest rate of 12%, which companys machine should Acme Engineering buy? (a) Use a present worth analysis to determine your recommendation to Acme. (b) Use an annual worth analysis to determine your recommendation to Acme.
Answer #1

ANSWER:

A) Present worth analysis:

Johnson punches :

initial charges = $150,000

maintenance charges = $4,000

savings = $89,000

salvage value = $15,000

i = 12%

n = 5 years

pw = initial charges + maintenance charges(p/a,i,n) + savings(p/a,i,n) + salvage value(p/f,i,n)

pw = -150,000 - 4,000(p/a,12%,5) + 89,000(p/a,12%,5) + 15,000(p/f,12%,5)

pw = -150,000 + 85,000(p/a,12%,5) + 15,000(p/f,12%,5)

pw = -150,000 + 85,000 * 3.605 + 15,000 * 0.5674

pw = -150,000 + 306,425 + 8,511

pw = 164,936

Riser machines:

initial charges = $205,000

maintenance charges = $4,300

savings = $86,000

salvage value = $15,000

i = 12%

n = 10 years

pw = initial charges + maintenance charges(p/a,i,n) + savings(p/a,i,n) + salvage value(p/f,i,n)

pw = -205,000 - 4,300(p/a,12%,10) + 86,000(p/a,12%,10) + 15,000(p/f,12%,10)

pw = -205,000 + 81,700(p/a,12%,10) + 15,000(p/f,12%,10)

pw = -205,000 + 81,700 * 5.65 + 15,000 * 0.3220

pw = -205,000 + 461,605 + 4,830

pw = 261,435

so based on present worth analysis riser machines should be chosen.

B) Annual worth analysis:

Johnson punches:

aw = pw(a/p,i,n)

i =12% , n = 5 years and pw = 164,936

aw = 164,936(a/p,12%,5)

aw = 164,936 * 0.2774

aw = 45,753.24

Riser machines:

aw = pw(a/p,i,n)

i =12% , n = 10 years and pw = 261,435

aw = 261,435(a/p,12%,10)

aw = 261,435 * 0.177

aw = 46,273.99

so based on annual worth analysis riser machines should be chosen.